A teacher's table is to be constructed such that the width is 3/5 of the length. If the perimeter is 72 inches find the area of the rectangle

Answer by Cromlix(4381)

Hi there,
Make length = x
Width = 3/5x
Perimeter = 2(length) + 2(width)
72 = 2(x) + 2(3/5x)
72 = 2x + 6/5x
72 = 16/5x
72*5/16 = x (* = times)
x = 22.5 inches
Length = 22.5 inches
Width = 13.5 inches
Area = length * width
Area = 22.5 * 13.5 = 303.75 ins^2
Hope this helps :-)